

********************************************************************
UARS WEEKLY STATUS REPORT - 09 SEP 2005
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 27 August 2005 (Orbit 76415, GMT Day 239) through
02 September 2005 (Orbit 76519, GMT Day 245).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 240/2003 76442 POWER ON
HALOE 241/0406 76447 POWER OFF
HALOE 241/2242 76458 POWER ON
HALOE 242/0716 76464 POWER OFF
HALOE 242/1118 76467 POWER ON
HALOE 242/1414 76469 POWER OFF
HALOE 243/0836 76480 POWER ON
HALOE 243/1631 76485 POWER OFF
SOLSTICE 243/1812 76486 POWER ON - DAY-ONLY
SOLSTICE 243/1952 76487 DAY-ONLY TO FULLTIME
The beta angle ranged between the angles of 65.9 degrees to 59.9
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 2 degrees on 17 September 2005 (DOY 260).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 3.80 to 3.15 degrees C. Additional
battery performance data for each battery on 26 August 2005 and 02
September 2005 are provided in Appendix D.
The clock error ranged from 11.8 to -11.9 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 16
SOLSTICE = 02
HALOE = 03
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 04
Five (5) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
GROUND SYSTEM OPERATION
Four (4) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT has been closely monitoring the condition of battery
2. Battery 2 continues its slow drop in temperature. Battery 3 has
been working surprisingly well under the increased load that it has
been put under. During the 241/1819 contact, our latest tweak to TMON
45 and 46 was made. TMON 45 was adjusted to take battery 2 off relay
once FCT + 1 count was reached. This was done to keep the C/D ratio
on battery 2 in check as it was increasing with the increase in S/C
beta angle. TMON 46 was also modified to keep battery 2 off relay
until 7 « minutes into S/C night. This change was necessary due to
battery 3 actually charging battery 2 when it came back on relay.
The FOT has also taken the first steps to begin orbit-lowering
activities. On the 242/1916 contact the FOT successfully executed the
FSW patch to allow for 4 thruster mode on the PM. In order to verify
that the PM would behave as expected with the new patch on-board, the
FOT had planned a short calibration burn of 6 seconds using all 4-
translation thrusters. This test was planned for and carried out with
out any problems on the 244/1330 support. All 4 thrusters fired as
expected and the A/C attitude thrusters also fired in a timely manner.
In anticipation of longer burns, 2 more calibration burns of
increasingly longer lengths are tentatively scheduled for 14 September
2005 (DOY 257) and 22 September 2005 (DOY 265). A schedule for when
the 6 required orbit-lowering burns has also been settled upon. Burn
1 will be on 29 September 2005 (DOY 272); burn 2 will be on 04 October
2005 (DOY 277); burn 3 will be on 06 October 2005 (DOY 279); burn 4
will be on 12 October 2005 (DOY 285); burn 5 will be on 18 October
2005 (DOY 291); and lastly burn 6 will be on 20 October 2005 (DOY
293). There will be no instruments operating during the month of
October due to the aggressive burn plan activities. HALOE will resume
operations after the burns have completed and we have completed the
yaw back to forward flight.
Lastly, the UARS spacecraft flight software rejected the ephemeris
that was created for 02 September 2005, (DOY 245). This was due to
the fact that the calibration burn that was done the day prior moved
UARS enough to require the FOT to open the ephemeris check window.
This small change was made and a new ephemeris was uploaded, accepted
and began propagating as expected. In the future, the FOT will have
to remember to open the windows for any ephemeris loads sent after any
burns.
MEETINGS
The FOT attended the De-boost status meeting on 29 August 2005.
During this meeting FDF proposed burns of 18 minutes in length for
each orbit lowering burn instead of the 15-minute burns that they
initially had suggested might be needed. These burns are based on a
duty cycle of ~ 90%.
The FOT also attended the weekly Technical Working Group meeting on 01
September 2005. FDF predicts that UARS fuel remaining onboard is in
the range of 354 lbs plus or minus 3 lbs. UARS telemetry shows there to be
353.5 lbs of fuel remaining.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 September 2005 (DOY 260).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
A second 4-thruster calibration burn of roughly 15 seconds in length
is scheduled for 14 September 2005 (DOY 257).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 02 September 2005
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
76433 240/0543 05-288 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
76439 240/1453 05-289 OBC/CLK CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
76442 240/2003 05-290 HALOE POWER ON - PART ONE
76444 240/2246 05-291 HALOE POWER ON - PART TWO
76447 241/0406 05-292 HALOE POWER OFF
76454 241/1527 05-293 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
76455 241/1637 05-294 OBC OBC TABLE DUMPS
76456 241/1711 05-295 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
76456 241/1819 05-296 TMON TMON 45/46 CHANGE
76457 241/1856 05-297 OBC OBC BANK DUMPS
76458 241/2242 05-290 HALOE POWER ON - PART ONE
76460 242/0032 05-291 HALOE POWER ON - PART TWO
76464 242/0716 05-292 HALOE POWER OFF
76467 242/1118 05-290 HALOE POWER ON - PART ONE
76468 242/1229 05-291 HALOE POWER ON - PART TWO
76469 242/1414 05-292 HALOE POWER OFF
76470 242/1554 05-298 OBC MEMORY MONITOR
76472 242/1916 05-300 OBC/PM PM PATCH (4-THRUSTER)
76472 242/1916 05-299 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
76473 242/2100 05-301 OBC OBC BANK DUMPS
76473 242/2135 05-302 OBC/CLK CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
76480 243/0836 05-290 HALOE POWER ON - PART 1
76481 243/1016 05-291 HALOE POWER ON - PART 2
76485 243/1631 05-293 HALOE POWER OFF
76486 243/1812 05-303 SOLSTICE POWER ON - DAY ONLY
76487 243/1952 05-304 SOLSTICE POWER ON - FULLTIME
76489 243/2239 05-305 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
76493 244/0430 05-306 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
76498 244/1330 05-307 PM OA CAL BURN
76517 245/1909 05-308 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 02 September 2005
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 02 September 2005
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
76433 240/0546 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
76439 240/1505 CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
76454 241/1530 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
76456 241/1720 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
76472 242/1924 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
76473 242/2137 CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
76489 243/2239 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
76493 244/0430 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
76517 245/1909 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 02 September 2005
26 August 2005 - GMT Day 238 Beta = 62.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -3.68/ -4.63 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 +7.74/ +3.87 +22.0/-17.6 25.8 1.09 11.8
3 196.0/-700.0 +3.87/ +2.60 +19.2/-20.8 25.6 1.03 16.3
02 September 2005 - GMT Day 245 Beta = 59.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 728.0/ 728.0 -2.74/ -3.99 +00.0/+00.0 04.6 0.00 0.0
2 -700.0/-700.0 +1.34/ +0.40 +16.8/-13.2 25.4 1.08 7.2
3 128.8/-291.2 +5.14/ +3.55 +27.2/-24.0 25.1 1.02 21.2
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 02 September 2005
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-043 FOT 76458 241/2148 40662 Late Acquisition
05-032 FOT 76465 242/0757 30832 Generic Late Acquisition
05-045 FOT 76496 244/0901 40682 Late Acquisition/Equip @
WSGT
05-046 OBC 76518 245/2125 N/A Flight Segment Ephem
Reject
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
05-042 MPS 76339 233/2347 N/A BATTERY 2 DIFFV -700mV
05-043 FOT 76458 241/2148 40662 Late Acquisition
05-032 FOT 76465 242/0757 30832 Generic Late Acquisition
05-045 FOT 76496 244/0901 40682 Late Acquisition/Equip @
WSGT
05-046 OBC 76518 245/2125 N/A Flight Segment Ephem
Reject